# MuLan: Multimodal-LLM Agent for Progressive Multi-Object Diffusion

### v1

@@ Abstract
Existing text-to-image models still struggle to generate images of multiple objects, especially in handling their spatial positions, relative sizes, overlapping, and attribute bindings. In this paper, we develop a training-free Multimodal-LLM agent (MuLan) to address these challenges by progressive multi-object generation with planning and feedback control, like a human painter. MuLan harnesses a large language model (LLM) to decompose a prompt to a sequence of sub-tasks, each generating only one object conditioned on previously generated objects by stable diffusion. Unlike existing LLM-grounded methods, MuLan only produces a high-level plan at the beginning while the exact size and location of each object are determined by an LLM and attention guidance upon each sub-task. Moreover, MuLan adopts a vision-language model (VLM) to provide feedback to the image generated in each sub-task and control the diffusion model to re-generate the image if it violates the original prompt. Hence, each model in every step of MuLan only needs to address an easy sub-task it is specialized for. We collect 200 prompts containing multi-objects with spatial relationships and attribute bindings from different benchmarks to evaluate MuLan. The results demonstrate the superiority of MuLan in generating multiple objects over baselines. The code is available on https:github.com/measure-infinity/mulan-code.
Computer Vision, Machine Learning, ICML
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) have shown growing potential in generative AI tasks, especially in creating diverse and high-quality images with text prompts (Saharia et al., 2022; Rombach et al., 2022). However, current state-of-the-art text-to-image (T2I) models such as Stable Diffusion (Rombach et al., 2022) and DALL-E 3 (Betker et al., 2023) still struggle to deal with complicated prompts involving multiple objects and lack precise control of their spatial relations, potential occlusions, relative sizes, etc. As shown in Figure 2, to generate a sketch of “The orange pumpkin is on the right side of the black door”, even the SOTA open-source T2I model, Stable Diffusion XL (Podell et al., 2023), still generates wrong attribute-binding as well as incorrect spatial positions of several objects.
Among works that aim to improve the controllability of T2I models on complicated prompts, a recent promising line of research seeks to utilize large language models (LLMs), e.g., ChatGPT, GPT-4 (Achiam et al., 2023), to guide the generation process (Lian et al., 2023; Feng et al., 2023). Specifically, an LLM is prompted to generate a layout for the given prompt, i.e., a bounding box for each object in the image, given detailed instructions or demonstrations if necessary. However, due to the limited spatial reasoning capability of LLMs as well as their lack of alignment with the diffusion models, it is still challenging for LLMs to directly generate a complete and precise layout for multiple objects. Without a feedback loop interacting with the generative process, the layout’s possible mistakes cannot be effectively detected and corrected. Moreover, the layout is often applied as an extra condition in addition to the original prompt (e.g., bounding boxes combined with GLIGEN (Li et al., 2023)), so the diffusion models may still generate an incorrect image due to its misunderstanding of the complicated prompt.
@@
To address the limitations and challenges of previous methods, we develop a training-free and controllable T2I generation paradigm that does not require demonstrations but mainly focuses on improving the tool usage of existing models. Our paradigm is built upon a progressive multi-object generation by a Multimodal-LLM agent (MuLan), which generates only one object per stage, conditioned on generated objects in the image and attention masks of the most plausible positions to place the new object. Unlike previous methods that add conditions to each model and make the task even more challenging, MuLan uses an LLM as a planner decomposing the original T2I task into a sequence of easier subtasks. Each subtask generates one single object, which can be easily handled by diffusion models. To be noted, the LLM applied at the beginning of MuLan only focuses on high-level planning rather than a precise layout of bounding boxes, while the exact size and position of each object are determined later in each stage by LLM and attention guidance based on the generated objects in the image. Hence, we can avoid mistakes in the planning stage and find a better placement for each object adaptive to the generated content and adhering to the original prompt. In addition, MuLan builds a feedback loop monitoring the generation process, which assesses the generated image per stage using a vision-language model (VLM). When the generated image violates the prompt, the VLM will adjust the diffusion model to re-generate the image so any mistake can be corrected before moving to the next stage. Furthermore, we develop a strategy applied in each stage to handle the overlapping between objects, which is commonly ignored by previous work (Lian et al., 2023).
@@
Therefore, MuLan obtains better controllability of the multi-object composition. An illustration of the progressive generation process is shown in Figure 1. To evaluate MuLan, we curate a dataset of intricate and challenging prompts from different benchmarks. To compare MuLan with existing approaches, we prompt GPT-4V (OpenAI, 2023) several questions based on the input texts to comprehensively evaluate the alignment of the generated images with the prompts from three aspects. We further conduct human evaluations of the generated images. Extensive experimental results show that MuLan can achieve better controllability over the generation process and generate high-quality images aligning better with the prompts than the baselines. Example images generated by different methods are shown in Figure 2. Our main contributions are summarized as follows:
@@
Figure 2: Examples of MuLan-generated images, compared to the original SD-v1.4 (Rombach et al., 2022), the original SDXL (Podell et al., 2023), Structure diffusion (Feng et al., 2022), Promptist (Hao et al., 2022), and PixArt-$\alpha$ (Chen et al., 2023).
@@ 2 Related Work
As a new family of generative models, diffusion models have attracting more and more attention due to its powerful creative capability. Text-to-image generation, which aims to generate the high-quality image aligning with given text prompts, is one of the most popular applications (Nichol et al., 2021; Saharia et al., 2022; Rombach et al., 2022; Betker et al., 2023). Among different powerful diffusion models, the latent diffusion model (Rombach et al., 2022) has shown amazing capability and has been widely used in practice due to the efficiency and superior performance, which is also the backbone of the current SOTA stable diffusion models. Different from the typical diffusion models which directly perform the diffusion and denoising process in the pixel space, the latent diffusion model perform the whole process in the encoded latent space (Rombach et al., 2022), which can greatly reduce the training and inference time. Recently, empowered by a significantly expanded model capacity, Stable Diffusion XL has demonstrated performance levels approaching commercial application standards (Podell et al., 2023).
@@
Although Stable Diffusion model has shown unprecedented performance on the T2I generation task, it still struggles with text prompts with multi-object, especially when there are several spatial relationships and attribute bindings in the prompts. To achieve more controllable and accurate image compositions, many compositional generation methods have been proposed. StructureDiffusion (Feng et al., 2022) proposed a training-free method to parse the input prompt and combine it with the cross-attention to achieve better control over attribute bindings and compositional generation. On the other hand, Promptist (Hao et al., 2022) aimed to train a language model with the objective of optimizing input prompts, rendering them more comprehensible and facilitative for diffusion models. Recently, several works utilize the large language model to directly generate the whole layout for the input prompt with in-context learning, and then generate the image conditioned on the layout (Lian et al., 2023; Feng et al., 2023). While all the previous take the whole input prompt, we propose to turn the original complicated task into several easier sub-tasks. A training-free multimodal-LLM agent is utilized to progressively generate objects with feedback control so that the whole generation process would be better controlled.
@@ 3 Multimodal-LLM Agent (MuLan)
3.1 Problem Formulation
Likewise, MuLan begins by strategically planning and decomposing an intricate input prompt into a manageable sequence of sub-prompts, each focusing on an easier sub-task generating one single object. MuLan then adopts a progressive strategy that generates one object in each stage conditioned on previously generated objects using a diffusion model. Simultaneously, a VLM offers insightful feedback and adaptively adjusts the generation process to guarantee precision in accomplishing each subtask. Compared to previous methods, MuLan is entirely training-free and does not require any examples. As illustrated in Fig. 1, MuLan is composed of three components:
@@ 3.1 Problem Formulation
  • Conditional single-object diffusion with LLM planning and attention guidance, which generates a new object conditioned on the previous step’s image using a stable diffusion model. While a sub-prompt from LLM planning provides text guidance, the object’s size and position are controlled by an attention mask.
  • VLM-feedback control, which inspects the image generated per stage and adjusts hyperparameters to re-generate the image if it violates the original prompt.
3.2 Background on (Latent) Diffusion Models
Consisting of the diffusion process and the reverse process, diffusion models have shown impressive capability for high-quality image generation by iteratively adding noise and denoising (Ho et al., 2020). Let $\bm{x}_{0}\sim q(\bm{x}_{0})$ be the true data distribution. Starting from $\bm{x}_{0}$, the diffusion process adds different levels of noise pre-defined by the schedule $\{\beta_{t}\}_{1}^{T}$, producing $\bm{x}_{1},\cdots,\bm{x}_{T}$. As $T\to\infty$, $\bm{x}_{T}$ will become the standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$. Accordingly, the reverse process aims to reverse the above process and reconstruct the true data distribution from $p(\bm{x}_{T})=\mathcal{N}(\bm{0},\bm{I})$ by a parameterized noise model $\epsilon_{\theta}(\cdot)$. With $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, the training loss of the model can be simplified as
Latent diffusion models (Rombach et al., 2022) have recently attracted growing attention due to their efficiency and superior performance. Instead of performing diffusion and its reverse process in the pixel space, they add noise and denoise in a latent space of $\bm{z}$ encoded by a pre-trained encoder $\mathcal{E}$. Thereby, the diffusion process starts from $\bm{z}_{0}=\mathcal{E}(\bm{x}_{0})$ and subsequently produces latent states $\bm{z}_{1},\cdots,\bm{z}_{t},\cdots,\bm{z}_{T}$. Accordingly, the training loss becomes
@@ 3.3 Prompt Decomposition by LLM Planning
Given a complicated prompt p, MuLan first uses an LLM to automatically decompose p into $N$ object-wise sub-prompts $\texttt{p}_{1:N}$. MuLan specifically asks the LLM to produce a sequence of objects that will be created in the default order from left to right and bottom to top in the image, so the LLM leverages its prior knowledge to fill all objects of p to an empty list of the pre-defined order. Compared to layout generation, this is more feasible for most LLMs who may suffer from hallucinations in spatial reasoning. The pre-defined order will also facilitate sequential object generation in Section 3.4. Let $\texttt{objs}=\{\texttt{obj}_{1},\cdots,\texttt{obj}_{n},\cdots,\texttt{obj}_{N}\}$ be the LLM-planned $N$ objects extracted from p. For the first object, the sub-prompt is simply $\texttt{p}_{1}=$$\{\texttt{obj}_{1}\}$”. For object-$n$ with $n>1$, the subtask is to generate object-$n$ conditioned on previous objects and the textual sub-prompt is defined as $\texttt{p}_{n}=$$\{\texttt{obj}_{n}\}\,\,\texttt{and}\,\,\{\texttt{obj}_{n-1}\}$”. MuLan conducts the above global planning by an LLM at the very beginning before generating any image. The detailed prompts and template for LLM planning can be found in Appendix A.

### v2

# MuLan: Multimodal-LLM Agent for Progressive and Interactive Multi-Object Diffusion

@@ Abstract
Existing text-to-image models still struggle to generate images of multiple objects, especially in handling their spatial positions, relative sizes, overlapping, and attribute bindings. To efficiently address these challenges, we develop a training-free Multimodal-LLM agent (MuLan), as a human painter, that can progressively generate multi-object with intricate planning and feedback control. MuLan harnesses a large language model (LLM) to decompose a prompt to a sequence of sub-tasks, each generating only one object by stable diffusion, conditioned on previously generated objects. Unlike existing LLM-grounded methods, MuLan only produces a high-level plan at the beginning while the exact size and location of each object are determined upon each sub-task by an LLM and attention guidance. Moreover, MuLan adopts a vision-language model (VLM) to provide feedback to the image generated in each sub-task and control the diffusion model to re-generate the image if it violates the original prompt. Hence, each model in every step of MuLan only needs to address an easy sub-task it is specialized for. The multi-step process also allows human users to monitor the generation process and make preferred changes at any intermediate step via text prompts, thereby improving the human-AI collaboration experience. We collect 200 prompts containing multi-objects with spatial relationships and attribute bindings from different benchmarks to evaluate MuLan. The results demonstrate the superiority of MuLan in generating multiple objects over baselines and its creativity when collaborating with human users. The code is available at https://github.com/measure-infinity/mulan-code.
@@ 1 Introduction
Diffusion models [19, 9, 20] have shown growing potential in generative AI tasks, especially in creating diverse and high-quality images with text prompts [18, 17]. However, current state-of-the-art text-to-image (T2I) models such as Stable Diffusion [17] and DALL-E 3 [2] still struggle to deal with complicated prompts involving multiple objects and lack precise control of their spatial relations, potential occlusions, relative sizes, etc. As shown in Figure 2, to generate a sketch of “The orange pumpkin is on the right side of the black door”, even the SOTA open-source T2I model, Stable Diffusion XL [16], still generates wrong attribute-binding as well as incorrect spatial positions of several objects.
Among works that aim to improve the controllability of T2I models on complicated prompts, a recent promising line of research seeks to utilize large language models (LLMs), e.g., ChatGPT, GPT-4 [1], to guide the generation process [12, 6]. Specifically, an LLM is prompted to generate a layout for the given prompt, i.e., a bounding box for each object in the image, given detailed instructions or demonstrations if necessary. However, due to the limited spatial reasoning capability of LLMs as well as their lack of alignment with the diffusion models, it is still challenging for LLMs to directly generate a complete and precise layout for multiple objects. Without a feedback loop interacting with the generative process, the layout’s possible mistakes cannot be effectively detected and corrected. Moreover, the layout is often applied as an extra condition in addition to the original prompt (e.g., bounding boxes combined with GLIGEN [11]), so the diffusion models may still generate an incorrect image due to its misunderstanding of the complicated prompt.
@@
To address the limitations and challenges of previous methods, we develop a training-free and controllable T2I generation paradigm that does not require demonstrations but mainly focuses on improving the tool usage of existing models. Our paradigm is built upon a progressive multi-object generation by a Multimodal-LLM agent (MuLan), which generates only one object per stage, conditioned on generated objects in the image and attention masks of the most plausible positions to place the new object. Unlike previous methods that add conditions to each model and make the task even more challenging, MuLan uses an LLM as a planner decomposing the original T2I task into a sequence of easier subtasks. Each subtask generates one single object, which can be easily handled by diffusion models. To be noted, the LLM applied at the beginning of MuLan only focuses on high-level planning rather than a precise layout of bounding boxes, while the exact size and position of each object are determined later in each stage by LLM and attention guidance based on the generated objects in the image. Hence, we can avoid mistakes in the planning stage and find a better placement for each object adaptive to the generated content and adhering to the original prompt. In addition, MuLan builds a feedback loop monitoring the generation process, which assesses the generated image per stage using a vision-language model (VLM). When the generated image violates the prompt, the VLM will adjust the diffusion model to re-generate the image so any mistake can be corrected before moving to the next stage. Furthermore, we develop a strategy applied in each stage to handle the overlapping between objects, which is commonly ignored by previous work [12].
@@
Therefore, MuLan obtains better controllability of the multi-object composition. An illustration of the progressive generation process is shown in Figure 1. Note that there is a concurrent work called RPG [22] sharing a similar high-level idea (i.e., decomposing the prompt into sub-tasks) with MuLan. However, there still exist substantial differences between ours and RPG. MuLan generates each object conditioned on previously generated objects while RPG generates all objects independently. MuLan does not require any manually designed demonstrations for in-context learning. In addition, as shown in Section 4.1, MuLan can be directly applied to human-agent interaction during generation, which greatly boosts the flexibility and effectiveness of the generation. To evaluate MuLan, we curate a dataset of intricate and challenging prompts from different benchmarks. To compare MuLan with existing approaches, we prompt GPT-4V [15] several questions based on the input texts to comprehensively evaluate the alignment of the generated images with the prompts from three aspects. We further conduct human evaluations of the generated images. Extensive experimental results show that MuLan can achieve better controllability over the generation process and generate high-quality images aligning better with the prompts than the baselines. Example images generated by different methods are shown in Figure 2. Our main contributions are summarized as follows:
@@
Figure 2: Examples of MuLan-generated images, compared to the original SD-v1.4 [17], the original SDXL [16], Structure diffusion [5], Promptist [7], and PixArt-$\alpha$ [3].
@@
  • We show that the proposed framework can be applied to human-agent interaction during generation. This enables users to effectively monitor and change/adjust the generation process during generation instead of waiting until all the generation is finished.
@@ 2 Related Work
As a new family of generative models, diffusion models have attracting more and more attention due to its powerful creative capability. Text-to-image generation, which aims to generate the high-quality image aligning with given text prompts, is one of the most popular applications [14, 18, 17, 2]. Among different powerful diffusion models, the latent diffusion model [17] has shown amazing capability and has been widely used in practice due to the efficiency and superior performance, which is also the backbone of the current SOTA stable diffusion models. Different from the typical diffusion models which directly perform the diffusion and denoising process in the pixel space, the latent diffusion model perform the whole process in the encoded latent space [17], which can greatly reduce the training and inference time. Recently, empowered by a significantly expanded model capacity, Stable Diffusion XL has demonstrated performance levels approaching commercial application standards [16]. Detailed background on the procedure of diffusion models is provided in Appendix C.
@@
Although Stable Diffusion model has shown unprecedented performance on the T2I generation task, it still struggles with text prompts with multi-object, especially when there are several spatial relationships and attribute bindings in the prompts. To achieve more controllable and accurate image compositions, many compositional generation methods have been proposed. StructureDiffusion [5] proposed a training-free method to parse the input prompt and combine it with the cross-attention to achieve better control over attribute bindings and compositional generation. On the other hand, Promptist [7] aimed to train a language model with the objective of optimizing input prompts, rendering them more comprehensible and facilitative for diffusion models. Several works utilize the large language model to directly generate the whole layout for the input prompt with in-context learning, and then generate the image conditioned on the layout [12, 6]. While all the previous take the whole input prompt, we propose to turn the original complicated task into several easier sub-tasks. A training-free multimodal-LLM agent is utilized to progressively generate objects with feedback control so that the whole generation process would be better controlled. Very recently, a concurrent work RPG [22] also proposed to utilize LLM agent to decompose the prompt into different subtasks. However, MuLan generates each object step by step and correct mistakes after each step rather than treating all subtasks independently and does not need a well-designed in-context learning demonstrations. We defer a more thorough discussion with RPG [22] in Appendix B.
@@ 3 Multimodal-LLM Agent (MuLan)
3.1 Overview
MuLan begins by strategically planning and decomposing an intricate input prompt into a manageable sequence of sub-prompts, each focusing on an easier sub-task generating one single object. MuLan then adopts a progressive strategy that generates one object in each stage conditioned on previously generated objects using a diffusion model. Simultaneously, a VLM offers insightful feedback and adaptively adjusts the generation process to guarantee precision in accomplishing each subtask. Compared to previous methods, MuLan is entirely training-free and does not require any in-context examples. As illustrated in Fig. 1, MuLan is composed of three components:
@@ 3.1 Overview
  • Conditional single-object diffusion with LLM planning and attention guidance, which generates a new object conditioned on the previous step’s image using a stable diffusion model. While a sub-prompt from LLM planning provides text guidance, the object’s size and position are controlled by an attention mask, which guides the object to be correctly positioned and generated.
  • Feedback control by interacting with VLM, which inspects the image generated per stage and adjusts hyperparameters and attention guidance to re-generate the image if it violates the original prompt.
@@ 3.2 Prompt Decomposition by LLM Planning
Given a complex prompt p, MuLan first uses an LLM to automatically decompose p into $N$ object-wise sub-prompts $\texttt{p}_{1:N}$. During decompostion, MuLan specifically asks the LLM to produce a sequence of objects that will be created in the default order from left to right and bottom to top in the image. The LLM can easily finish this task by leveraging its prior knowledge to fill all objects of p to an empty list of the pre-defined order without in-context learning which requires manually designed examples. Let $\texttt{objs}=\{\texttt{obj}_{1},\cdots,\texttt{obj}_{n},\cdots,\texttt{obj}_{N}\}$ be the LLM-planned $N$ objects extracted from p. For the first object, the sub-prompt is simply $\texttt{p}_{1}=$$\{\texttt{obj}_{1}\}$”. For object-$n$ with $n>1$, the subtask is to generate object-$n$ conditioned on previous objects and the textual sub-prompt is defined as $\texttt{p}_{n}=$$\{\texttt{obj}_{n}\}\,\,\texttt{and}\,\,\{\texttt{obj}_{n-1}\}$”. MuLan conducts the above global planning by an LLM at the very beginning before generating any image. The detailed prompts and template for LLM planning can be found in Appendix E.
@@ 3.3 Conditional Single-Object Diffusion with LLM Planning and Attention Guidance
At stage-$n$, the diffusion model only focuses on generating $\texttt{obj}_{n}$ according to the sub-prompt $\texttt{p}_{n}$, ensuring that $\texttt{obj}_{n}$ can be correctly positioned and generated. To this end, MuLan utilizes the LLM to plan the relative position and size of $\texttt{obj}_{n}$, allocating a rough mask (i.e., a bounding box) $\bm{M}_{n}$ for $\texttt{obj}_{n}$. Then, cross-attention guidance is applied during the generation of $\texttt{obj}_{n}$ to ensure $\texttt{obj}_{n}$ is appropriately positioned within $\bm{M}_{n}$. The pipeline is given in Figure 3 with the complete procedure listed in Algorithm 1 in Appendix D. We will introduce it step by step in the following.